OpenStudy (anonymous):

How would I graph this function: y=x^2/3

OpenStudy (anonymous):

Since I know that any number I use as my test point is to be evaluated using this formula x^a/b = b^(sqrt x^a) where a is 3 and b is 2, should I only choose values that can be simplified with the cubed sqrt like 8, and 27 and not a number like 1 or 0?

OpenStudy (anonymous):

1 and 0 can be simplified with the cubed sqrt, so I'd use them.

OpenStudy (anonymous):

But you need more points than that.

OpenStudy (anonymous):

Can you graph y=x^2?

OpenStudy (anonymous):

yeah, that would just be a parabola.

OpenStudy (anonymous):

If \(y=x^2\) is like this,.Then, \(y=\frac{1}{3}x^3\) must be like (Drawing)

OpenStudy (anonymous):